3. (ICs for Quasi-Linear Preferences) Consider the utility function: u(x, y) = x1/2 + y.

a. Find the expression for the MRS (= – dy/dx).

b. Draw one IC making sure its shape reflects your expression for MRS above.

c. Given your expression for MRS, draw another IC above the one you just drew, and comment on how the slopes of the ICs compare at a given level of x (e.g., at x = 1).
Answer #1

a) MRS= MUx/MUy

MUx= 1/2√x

MUy=1

MRS= 1/2√x

MRS=0.5/√x

b)

C)The slope of IC remains the same at a given level of x on different ICs.

When x=1, MRS= 1/2√1= 1/2